OpenStudy (anonymous):

Find the derivative of f(x) = -2x^2 + 11x at x=9

OpenStudy (anonymous):

can someone help me solve this while we use the formula \[\lim_{h \rightarrow 0} \frac{ f(x+h)-f(x) }{ h } \]

OpenStudy (anonymous):

I somehow do not know how to set it up properly

OpenStudy (anonymous):

can you help me @Loser66

OpenStudy (loser66):

f(x+h)=??

OpenStudy (anonymous):

f(x+h) = -2(x+h)^2+11x?

OpenStudy (loser66):

\(f(x+h)= -2(x+h)^2 +11(x+h)\)

OpenStudy (anonymous):

why is there another (x+h)? what about the -f(x) part?

OpenStudy (loser66):

because f(x) = -2x^2 +11x hence, when replace x by x +h, you MUST replace all x on the expression

OpenStudy (anonymous):

okay, so instead of x its (x+h), but in the formula are we still going to use the other top half -f(x)?

OpenStudy (anonymous):

would it be -2(x+h)^2+11(x+h) - (-2x^2+11x)/ h?

OpenStudy (loser66):

yup, open parentheses, simplify . What do you have?

OpenStudy (anonymous):

okay let me see

OpenStudy (anonymous):

-4hx - 2h^2 + 11h?

OpenStudy (loser66):

you forgot h at the denominator. That is just numerator. From it, factor h out, cancel out with h in denominator, what do you have?

OpenStudy (anonymous):

yeah i meant just the numerator lol

OpenStudy (anonymous):

if i cancel the h, i would have -4x - 2h + 11

OpenStudy (anonymous):

would i now replace h with 0?

OpenStudy (loser66):

Hold on, your calculation is not correct. Let check

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

would i take -h out and have 4x + 2h - 11?

OpenStudy (loser66):

oh, you are correct

OpenStudy (loser66):

no need to take - out, just h

OpenStudy (anonymous):

okay i was confused for a second

OpenStudy (loser66):

cancel out, and get -4x -2h +11. Now take lim when h --> 0 we have -4x +11 , replace x =9. Dat sit

OpenStudy (anonymous):

-25

OpenStudy (loser66):

yup

OpenStudy (anonymous):

that should be it! I get it better now! Thank you very very much!

OpenStudy (loser66):

np